# MCVO: A Generic Visual Odometry for Arbitrarily Arranged Multi-Cameras

### v1

@@ III-E Multi-Camera Loop Closure
Although the multi-camera VO and scale correction provide high-accuracy state estimation, the accumulated drifts still exist alongside the running of pure visual odometry. To mitigate the drifts over time, we design a multi-camera loop closure module across different cameras and establish loop connections in the pose graph for further state optimization.
Loop Detection. In our MCVO pipeline, the loop detection simply leverages DBoW2 [33] algorithm for place recognition. The key point is how to use multi-camera features to improve the loop detection accuracy and robustness. Firstly, the features from all cameras are stored in the corresponding bag-of-words model. Then potential loop frames are identified by comparing matching scores between DBoW2 models from different frames. The wide field of view offered by multi-camera setups improve the detection rate of potential loops.
@@
Upon detecting a loop, we proceed with re-localization as illustrated in Fig. 4. The current frame features are converted into 3D features $\bm{z}$ with their depth information. By matching feature descriptors from the looped frames, we establish 2D-3D feature correspondences $\{(\bm{x}_{\textit{i}},\bm{z}_{\textit{i}})\}$. When a camera has sufficient 2D-3D feature pairs, we proceed with the following optimization to refine camera poses
$\min_{\mathcal{R},\mathcal{P}}\sum_{\mathcal{C}}\sum_{i}\big{|}\big{|}\bm{x}_{\textit{i}}-\bm{\pi}(\bm{K}_{\textit{c}},\bm{z}_{\textit{i}},\bm{R},\bm{p})\big{|}\big{|}^{2},$ (4)
where $\bm{K}_{\textit{c}}$ is the intrinsic and the extrinsic (to body frame) parameters of camera $c$, $\bm{\pi}$ is the projection of the $\bm{z}_{\textit{i}}$ on the image plane. The least squares problem is also solved by the Levenberg-Marquardt algorithm.
This process ensures accurate re-localization even with significant viewpoint differences. If multiple cameras meet the criteria, we use extrinsic constraints to verify the relocalization’s success further. Once verified, the accurate body pose of the loop frame is determined based on the camera poses and extrinsic parameters. Compared to a monocular system, our multi-camera setup enhances loop detection from different viewpoints, significantly increasing the detection rate of loops.
@@
VINS-Fusion 76.198 114.923 1.003 38.213 125.079 0.585 14.749 123.580 0.181 135.536 130.248 2.924 16.579 10.921% 14.80 16.469%
ORBSLAM3 22.673 117.981 0.390 12.269 115.832 0.306 8.134 114.361 0.164 44.317 114.268 1.484 2.90 1.91% 0.85 0.80%
MultiCamSLAM 15.68(245.2) 115.516 0.341 8.91(222.5) 114.082 0.184 11.28(91.9) 155.944 0.150 56.46(553.5) 136.175 1.561 14.83 9.75% 1.69 1.88%
MCVO(ORB) 44.054 8.634 0.518 64.861 4.803 1.367 49.143 5.353 0.591 102.050 6.139 2.833 11.50 7.561% 14.58 16.321%
MCVO(ours) 13.933 6.847 0.300 20.234 2.585 0.394 6.895 3.341 0.079 76.282 7.348 2.638 7.894 5.200% 1.54 1.711%
@@
Pose Graph Optimization We follow the commonly used 4-DoF pose graph optimization [1] to update the trajectory in the loops. The pose graph $\bm{G}=\{\bm{V},\bm{E}\}$ keep track of all keyframes with their spatial constraints, where vertices $\bm{V}$ represent keyframes and edges $\bm{E}$ store the transformation constraints between keyframe poses. Keyframes marginalized during the sliding window are added to the pose graph, with their relative transformations to other keyframes in the local sliding window incorporated as edges. When a loop is detected, a loop edge between the current frame and the loop frame is added to $\bm{E}$, followed by the 4-DoF pose graph optimization. The optimization process effectively reduces accumulated errors and enhances the system’s robustness.
@@ IV Experiments
In this section, we conduct experiments on two public datasets to evaluate our proposed method.
@@ IV-A1 Datasets
For the evaluation of our method, the dataset must contain multi-camera setup with pre-calibrated intrinsic and extrinsic parameters. We select two public available dataset, i.e., KITTI360[5] and MultiCamData[7], which have different camera types, scenario types, and various camera layouts. The KITTI360 dataset was collected in outdoor road environments with two $180^{\circ}$ fisheye cameras and two $90^{\circ}$ pinhole cameras, as shown in Fig. 1. The four cameras are synchronized and pre-calibrated. Compared to other datasets such as EuRoC [34], KITTI360 features larger scale, longer distances, and challenging scenarios, such as roads under bridges (lighting changes), wilderness (texture degradation), and dynamic scenes. We tested on most of the sequences and compared with the ground truth trajectories. MultiCamData was collected in indoor environments using 6 cameras. We mainly use the non-overlapping cameras (cam0, cam5, and cam6) for generalizability. The dataset also includes challenging scenarios, such as narrow corridors, large white walls (texture degradation), sudden turns, and dynamic scenes.
@@ IV-A2 Evaluation Setup
We select VINS-Fusion [1], ORB-SLAM3 [2], and MultiCamSLAM [7] as baselines for comparison. To obtain metric scale state estimation for VINS-Fusion and ORB-SLAM3, these two methods are setup with front stereo cameras of each dataset. While for MultiCamSLAM, we employ a non-overlapping camera setup consistent with the method [7]. MCVO is configured with two fisheye cameras and one front pinhole camera on KITTI360 dataset. The evaluation metrics on KITTI360 include Absolute Trajectory Error (ATE), Relative Pose Error (RPE) and scale error [35]. Since MultiCamData lacks ground truth trajectory, we estimate the errors between the initial and final poses using a visual target as in [7], given that the start and end of the trajectories are at the same location. All the experiments are conducted on a desktop with a 2.5GHz Intel Core i7-11700 CPU and an NVIDIA RTX3060 GPU.
@@ IV-B Experimental Results
The quantitative results for ATE are presented in Table I and Fig. 5. Compared with the stereo ORB-SLAM3 and VINS-Fusion, MCVO achieves significantly lower rotation error and competitive translation error on most of the sequences of KITTI360, primarily due to the wide FoV provided by the multi-camera setup and the high accuracy and robustness of the frontend. The small scale errors further demonstrate the effectiveness of our scale estimation strategy. The MultiCamSLAM method cannot finish most of the sequences on KITTI360. This issue arises because the KITTI360 dataset was captured at high speeds, leading to large inter-frame displacements. MultiCamSLAM struggles with significant inter-frame displacements, hindering triangulation and initialization. The comparison with MCVO on ORB features further shows the robustness of our front-end design. Besides, MCVO shows strong performance and robustness on MultiCamData. It outperforms VINS-Fusion in accuracy, with a marginal shortfall compared to ORB-SLAM3. This slight reduction in accuracy is primarily attributed to the presence of numerous white walls in indoor corridors, which introduce noise that impacts multi-camera VO. Compared to MultiCamSLAM, which also employs a non-overlapping camera setup, MCVO achieves superior accuracy and better generalization across different sequences.
Fig. 6 presents the RPE results of different methods on the 00 sequence of KITTI360 dataset. MCVO significantly outperforms other methods with smallest pose drift. This improvement is largely due to the wide coverage of scenes in the multi-camera setup with a robust and efficient frontend. The expanded FoV offers stronger constraints for state estimation, resulting in higher pose estimation accuracy. Additionally, the robust frontend ensures accurate feature tracking even in low-quality scenes. These results validate the effectiveness of MCVO on state estimation.
Figure 6: Comparison of RPE for different methods on the KITTI360_00 sequence.
To further qualitative analyze the performance, we plot the trajectories of different methods on sequences 00 and 05 of KITTI360 dataset in Fig. 7. ORB-SLAM3 and VINS-Fusion exhibit large rotation errors on the 00 sequence. In contrast, MCVO shows improved stability and accuracy under such conditions. Additionally, the overall trajectory of our approach is smoother with more stable position estimation.
@@ IV-C Ablation Study
In this section, we conduct ablation experiments for different numbers of cameras, feature extractors, feature matching, and whether to use loop closure. The effectiveness of each part of our proposed method is verified through comprehensive analyses.
TABLE II: MCVO performance under different number of cameras on the KITTI360 00 sequence.
Method Transl.[m] Rot.[°] UsageCPU[%]
Mean Std Mean Std Mean
2 cams 45.523 43.042 8.180 7.263 203.2
3 cams 13.933 13.640 6.847 6.806 316.7
4 cams 20.673 22.262 7.671 7.812 443.6
IV-C1 Camera number
Table II shows the state estimation accuracy of MCVO under different number of cameras. MCVO on two cameras obtains relatively lower accuracy due to the limited FoV, which provides insufficient information. When the number of cameras increases to 3, the estimation accuracy reaches its peak. However, with 4 cameras, there is a slight decrease in accuracy. This is primarily because two fisheye and one perspective cameras have already covered enough surrounding information, the added new camera will bring more observation noises and computational burden for the system. Consequently, the gain from the new camera cannot contribute to improving accuracy.
TABLE III: MCVO performance under different feature extractors on the KITTI360 00 sequence.
Method Transl.[m] Rot.[°] UsageCPU[%] FPS
Mean Std Mean Std Mean
Shi-Tomasi 39.390 40.612 9.256 8.290 398.3 18.75
ORB 44.054 39.900 8.634 8.327 942.4 12.74
SuperPoint 13.933 13.640 6.847 6.806 316.7 19.50
IV-C2 Feature extractor
We use pose accuracy and CPU usage metrics to compare different feature extraction methods. Table III shows that SuperPoint achieves the highest accuracy compared with commonly used Shi-Tomasi and ORB. This is mainly because SuperPoint can extract multi-level features and is more robust to environmental changes. Additionally, the CPU usage of MCVO(SuperPoint) is significantly lower. This reduction in CPU usage is because of our GPU-accelerated frontend. Consequently, The shift of CPU usage also ensures the computation resource for back-end optimization of accurate pose estimation and efficiency.
IV-C3 Feature matching
Table IV shows the performance of MCVO on different feature-matching methods. Flow-based methods exhibit lower CPU usage and more inliers compared to matching-based methods (KNN, SuperGlue). Additionally, we count the mean tracking times of each feature. LK shows the longest tracking times compared to others, which provides stable landmarks for the backend pose estimation. The small number of track times of other methods leads to the pose initialization failures.
TABLE IV: MCVO performance under different feature matching methods (Superglue-SG).
@@ IV-C4 Loop closure
We also conduct experiments on the KITTI360 05 sequence to validate the effectiveness of the loop closure. The qualitative results are shown in Fig. 8. We can observe that the multi-camera loop closure effectively performs loop detection and closure, providing stable global constraints for back-end optimization. MCVO successfully corrects the cumulative error in the map, achieving more precise trajectory estimation.
@@ IV-D Discussion
  • Multi-camera configurations can improve VO robustness, but inevitably introduce new computational requirements. Our approach uses learning-based feature extraction to shift CPU consumption but introduces the need for GPU resources. The best approach may be to develop a specialized camera with feature association embedded in the sensor hardware.
  • MCVO requires the extrinsic parameters of arbitrarily arranged multi-cameras. When there is no overlap between FoV, the extrinsic calibration will be a challenging problem.
@@ V Conclusions and Future Work
In this paper, we propose a generic multi-camera visual odometry system, which has high flexibility on the camera setup with different camera types, with or without FoV overlap. The arbitrary arrangement of multi-cameras, not only improves the flexibility of the SLAM system but also achieves robust camera-only true-scale pose estimation. The designed learning-based feature association front-end utilizes GPU to effectively release the computational pressure of CPUs on multi-camera data processing and improves the stability of feature matching. The system functionalities of multi-camera features for scale initialization, back-end optimization, and loop closure are fully validated during the experiments. Comparisons with the state-of-the-art further demonstrate the effectiveness and robustness of MCVO on state estimation. Future work will focus on the deployment of MCVO on more platforms.

### v2

@@ III-E Multi-Camera Loop Closure
Despite high-accuracy multi-camera VO and scale correction, drift accumulation persists. To address the problem, we design a cross-camera loop closure module.
Loop Detection. We construct a cross-camera feature fusion mechanism based on the DBow2[32] model. By integrating synchronously captured features from multi-camera into a unified feature space and enforcing spatiotemporal consistency constraints for inter-frame feature matching, we establish cross-camera shared bag-of-words representations. Specifically, as shown in Fig. 4. This methodology effectively mitigates geometric constraint degradation caused by viewpoint disparities, significantly enhancing the robustness of loop detection in complex multi-camera environments.
@@
VINS-Fusion 76.188 30.918 1.008 38.213 37.502 0.556 14.516 30.054 0.181 39.235 33.082 1.464 16.579 10.921% 14.80 16.469%
ORB-SLAM3 22.433 11.262 0.396 12.744 6.555 0.346 8.320 5.868 0.171 34.842 8.477 1.336 2.90 1.91% 0.85 0.80%
MultiCamSLAM 15.307(245.2) 16.712 0.330 8.716(222.5) 10.382 0.181 6.744(91.9) 55.356 0.119 56.413(553.5) 43.253 1.568 14.83 9.75% 1.69 1.88%
MCVO(ORB) 8.613 8.241 0.134 12.690 4.542 0.368 11.986 3.646 0.145 36.316 5.812 1.503 11.50 7.561% 14.58 16.321%
MCVO(2-cam) 6.059 8.459 0.105 11.829 5.179 0.272 6.215 7.274 0.062 24.826 7.947 1.007 5.85 3.85% 4.35 4.83%
MCVO(3-cam) 13.933 6.847 0.300 20.234 2.585 0.394 6.270 2.578 0.065 36.599 4.212 1.262 7.89 5.20% 1.54 1.711%
MCVO(4-cam) 7.610 8.303 0.129 4.889 2.256 0.138 3.469 3.264 0.037 22.262 5.418 0.876 5.50 3.62% 0.95 1.06%
@@
Pose Graph Optimization We employ 6-DoF pose graph $\bm{G}=\{\bm{V},\bm{E}\}$ optimization [1] to refine trajectories using vertices $\bm{V}$ (keyframes) and edges $\bm{E}$ (transformations). Marginalized keyframes from the sliding window are incorporated into $\bm{G}$ with relative transformations as edges. Upon loop detection, a loop closure edge triggers 6-DoF optimization, mitigating accumulated errors and enhancing robustness through spatiotemporal constraint propagation.
@@ IV-A1 Datasets
For the evaluation of our method, we require datasets containing multi-camera setups with pre-calibrated intrinsic and extrinsic parameters. We select two publicly available benchmarks: KITTI-360 [5] and MultiCamData [7], which exhibit complementary characteristics in terms of camera configurations, environmental diversity, and operational challenges. The outdoor-oriented KITTI-360 dataset comprises 4 cameras with heterogeneous configurations: two $180^{\circ}$ fisheye and two $90^{\circ}$ pinhole cameras arranged on a vehicle (Fig. 1). This dataset provides large-scale urban road scenarios with three principal challenges: (1) Illumination variations, (2) Texture degradation in wilderness, and (3) Dynamic objects. We evaluated on 4 sequences for rigorous trajectory comparison with ground truth. Besides, the indoor MultiCamData features 6 pinhole cameras comprising four forward and two side cameras. It presents distinctive challenges including (1) Feature scarcity in narrow corridors, and (2) Severe texture degradation on blank walls. The highly constrained environments enable effective verification of the robustness and generalization ability of the proposed method.
@@ IV-A2 Evaluation Setup
We evaluate against three state-of-the-art visual odometry systems: VINS-Fusion [1], ORB-SLAM3 [2], and MultiCamSLAM [7]. For metric consistency, VINS-Fusion and ORB-SLAM3 utilize calibrated stereo configurations, while MultiCamSLAM follows its native multi-camera implementation[10]. Our method adopts four-camera fusion on KITTI-360: two fisheye + two pinhole cameras, with ablative configurations (2-cam: dual pinhole; 3-cam: dual fisheye + mono pinhole). Evaluation metrics comprise Absolute Trajectory Error (ATE), Relative Pose Error (RPE), and scale drift[33] for KITTI-360, while MultiCamData evaluate metrics via visual marker-based start-end pose discrepancy [7] due to the absence of ground truth trajectories. All experiments are conducted on a desktop with an Intel Core i7-11700 CPU and an NVIDIA RTX3060 GPU.
@@ IV-B Experimental Results
The quantitative results for the ATE metric are shown in Table I and Fig. 5. Stereo VINS-Fusion shows much worse results than stereo ORB-SLAM3. Compared with the stereo ORB-SLAM3 and VINS-Fusion, MCVO achieves lower rotation and translation error on most of the sequences of KITTI360. This improvement stems from two key advantages: (1) the robust scale estimation module reduces scale drift across all sequences, and (2) the multi-camera configuration enables stable feature tracking through wide FoV coverage. MultiCamSLAM fails to complete most sequences in KITTI360 due to inadequate handling of large inter-frame displacements in high-speed scenarios, preventing reliable triangulation and initialization.
Experimental results of MCVO with varying camera configurations are presented in the last three rows of Table I. We first compare MCVO under 4 camera setup using ORB feature and the proposed learning-based front-end. The comparison shows the robustness of our front-end design. In terms of translation and scale estimation, 3-cam setup operates in a fully non-overlapping configuration, significantly outperforming ORB-SLAM3 and VINS-Fusion with stereo setup, which can be attributed to our robust multi-camera scale estimation strategy. Both 2-cam and 4-cam setups employ forward stereo cameras, where their overlapping FoV contributes to enhanced translation accuracy. Notably, the 4-cam setup generally achieves superior performance compared to the 2-cam setup, demonstrating the effectiveness of multi-camera configurations in improving state estimation accuracy. Regarding rotation estimation, multi-camera configurations significantly outperform stereo setups, as their distributed FoV maintains feature observability during rapid rotations, extending tracking continuity by 3-5 frames.
Moreover, MCVO demonstrates consistent robustness on MultiCamData, notably outperforming VINS-Fusion in pose estimation accuracy while being competitive to ORB-SLAM3. This minor precision gap stems from texture-deprived environments like uniform white walls, that degrade feature matching fidelity. Compared to MultiCamSLAM, which also employs a non-overlapping camera setup, MCVO achieves superior accuracy and better generalization across different sequences.
Fig. 6 presents the RPE results of different methods on the 05 sequence of KITTI360. MCVO (4-cam) outperforms other methods with smallest pose drifts. This improvement is largely due to our continuous correction strategy and multi-camera wide FoV. These results validate the effectiveness of MCVO on state estimation. To further qualitatively analyze the performance, we plot the trajectories of different methods on sequences 00 and 05 of KITTI360 in Fig. 7. ORB-SLAM3 and VINS-Fusion exhibit large rotation errors on the 00 sequence. In contrast, MCVO shows improved stability and accuracy under the same conditions.
Figure 6: Comparison of RPE for different methods on the KITTI360_05 sequence.
@@ IV-C Ablation Study
In this subsection, we conduct ablation study on the front-end and back-end of MCVO to verify their effectiveness.
TABLE II: MCVO performance under different feature extractors on the KITTI360 00 sequence.
Method Transl.[m] Rot.[°] UsageCPU[%] UsageGPU[%] FPS
Mean Std Mean Std
Shi-Tomasi(CPU) 9.413 9.448 8.570 8.559 534.68 - 18.75
ORB(CPU) 8.613 8.595 8.241 8.217 942.4 - 12.74
Superpoint(CPU) 28.841 26.653 9.553 9.108 1023.6 - 9.02
SuperPoint(GPU) 7.610 7.793 8.303 8.445 495.42 43.75 19.50
TABLE III: MCVO performance under different feature matching methods (Superglue-SG).
@@
TABLE IV: MCVO performance under different feature selection on the KITTI360 00 sequence. (Spatial Feature Distribution-SFD)
Method SFD $\downarrow$ Transl.[m] Rot.[°]
Mean Std Mean Std
score-based 24.25 8.740 8.678 8.617 8.515
quad-tree 22.18 8.432 8.437 8.680 8.633
3-priority 21.95 7.610 7.793 8.303 8.445
IV-C1 Feature extractor
Experimental evaluations comparing feature extractors across pose accuracy and computational efficiency metrics reveal SuperPoint’s superior performance (Table II). It achieves a 18.4% higher pose accuracy than conventional Shi-Tomasi and ORB, attributable to its multi-scale feature extraction capability and enhanced robustness to environmental variations. Furthermore, MCVO (SuperPoint) reduces CPU utilization by 47.4% compared to ORB, enabled by the GPU-accelerated frontend architecture. This efficiency gain facilitates resource allocation for backend optimization processes, ensuring both high-precision pose estimation and real-time operational efficiency.
IV-C2 Feature matching
As shown in Table III, feature matching based on LK flow in MCVO demonstrates superior computational efficiency and robustness compared to descriptor approaches. Quantitative analysis of feature tracking duration reveals that the LK flow maintains the longest feature lifespan (avg. 16.18 frames), providing stable landmark constraints for backend optimization. In contrast, other methods exhibit short tracking lifespans, resulting in a high failure rate during pose initialization.
IV-C3 Feature selection
As shown in Table IV, we conducted ablation studies on the feature selection algorithm. We introduce a new metric, SFD, to quantify Spatial Feature Distribution by computing the feature density on uniform grids of images and derive the variance across all cells. Compared with score-based methods, our 3-priority algorithm demonstrates 9.48% higher feature dispersion, effectively alleviating the issue of excessive feature concentration in texture-rich regions. The proposed method also achieves 11.0% improvement in pose estimation accuracy over both quadtree and score-based algorithms, validating its effectiveness.
@@ IV-C4 Loop closure
To validate the efficacy of the multi-camera loop closure, we conducted systematic evaluations on the KITTI360 05 sequence. Qualitative analysis (Fig. 8) demonstrates that our multi-camera closure framework enables robust loop detection and closure, establishing reliable global constraints for backend optimization. The proposed MCVO effectively mitigates cumulative mapping errors through this mechanism, achieving enhanced state estimation accuracy.
@@ IV-D Discussion
  • Multi-camera configurations can improve VO robustness, but inevitably introduce new computational requirements. Our approach uses learning-based feature extraction to shift CPU consumption but introduces the need of GPU resources. The best approach may be to develop a specialized camera with feature association embedded in the sensor hardware.
  • MCVO requires the extrinsic parameters of arbitrarily arranged multi-cameras. When there is no FoV overlap between cameras, precise extrinsic calibration will also be a challenging problem.
@@ V Conclusions and Future Work
In this paper, we propose a robust multi-camera visual odometry system, which has high flexibility on the camera setup with different camera types, with or without FoV overlap. The arbitrary arrangement of multi-cameras improves the flexibility of the SLAM system on more real-world platforms. Through establishing multi-camera manifold consistency constraints on motion trajectories, we effectively calculate metric-scale poses during initialization and continuously refine metric-scale in the backend optimization. Additionally, we further boost the accuracy and efficiency of the pose estimation through the learning-based feature association frontend and the multi-camera loop closure. Comparisons with the state-of-the-arts demonstrate the effectiveness and robustness of MCVO on state estimation. Future work will concentrate on deploying MCVO across more platforms and in diverse scenarios while simultaneously developing dense depth reconstruction using multi-cameras.